If 30.0 mL of 5.0 M HCl is used to make a 500.0 mL aqueous solution, what is the molarity of the dilute solution?
1
Expert's answer
2020-10-08T13:58:32-0400

V1=30mL=0.03L

C1(HCl)=5 M

V2=500mL=0.5L

C2(HCl)-?


n(HCl)= V1*C1(HCl)=0.03*5=0.15(mol)

C2(HCl)=n(HCl)/V2=0.15/0.5=3 M

Answer: 3 M